Write the equation of the line through the point (5,2) with slope -7/-3.
A. in point slope form
B. in slope-intercept form
C. in standard form

Answer by lwsshak3(11628)

Write the equation of the line through the point (5,2) with slope -7/-3.
A. in point slope form
(y1-y2)=m(x1-x2)
(y-2)=(7/3)(x-5)
3y-6=7x-35
3y=7x-29
y=7x/3-29/3
..
B. in slope-intercept form
y=mx+b
y=7x/3+b
solve for b using coordinates of given point (5,2)
2=7*5/3+b
b=2-35/3=6/3-35/3=-29/3
y=7x/3-29/3
..
C. in standard form
Ax+By=C
7x-3y=29